OpenStudy (anonymous):

how long will it take P dollars to triple itself at 5% simple interest?

OpenStudy (anonymous):

hi ab do you know the formula for simple interest?

OpenStudy (anonymous):

I=Prt P=principal,r=rate,t=time in years,I=interest if you want to triple you principal at 5%=0.05 (3P)=P(.05)t t=(3P)/(.05P)=3/0.05=60years